# EPW varieties as moduli spaces on ordinary GM surfaces and special GM threefolds

### v1

@@ Abstract.
We show that the double dual EPW sextic and double EPW sextic associated with a strongly smooth Gushel–Mukai surface $S$ can be realized as moduli spaces of semistable objects with respect to a stability condition on the bounded derived category of $S$. Also, we observe that double dual EPW surface and double EPW surface associated with a special Gushel–Mukai threefold $X$ can be realized as moduli spaces of semistable objects on the Kuznetsov component of $X$. As an application, we refine a statement of Bayer and Perry about Gushel–Mukai threefolds with equivalent Kuznetsov components for the special ones.
@@ 1.3. Main results
The moduli space $\textsf{{M}}_{\sigma_{S}}(1,0,-1)$ of semistable objects with Mukai vector (1,0,1)(1,0,-1) is isomorphic to the double dual EPW sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ associated $S$.
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(2,-H,2)$ of semistable objects with Mukai vector (2,H,2)(2,-H,2) is isomorphic to the double EPW sextics $\tilde{Y}^{\geq 1}_{A(S)}$ associated $S$.
@@ 5.4. The flopping walls and identification
It remains to show that $Y_{S}\cong X_{S}$ contracts to the moduli space $\textsf{{M}}_{\sigma_{S}}(1,0,-1)$ through the same symplectic resolution $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A^{\perp}}^{\geq 1}$ as in Theorem 5.1.
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ is isomorphic to $Y^{\geq 1}_{A^{\perp}_{S}}$.
@@
where $Z\subset S$ is a length $2$ subscheme spanning a line not in $S$. Indeed, such a short exact sequence follows from the fact that the ideal sheaf $\mathscr{I}_{L/M_{S}}$ of every line $L\subset M_{S}$ can be written as the cokernel of a unique map $\mathscr{U}_{G}|_{M_{S}}\rightarrow\mathscr{V}_{G}|_{M_{S}}$, and each such map has the ideal sheaf of a line as the cokernel. One just takes the restriction onto $S\subset M_{S}$. The line spanned by $Z$ is not in $S$ because such a length two subscheme $Z$ has been already flopped in Proposition 5.13. So one contracts the Fano variety $F(M_{S})$ of lines, and therefore recovers $f^{+}\colon Y_{S}\rightarrow Y_{A(S)^{\perp}}^{\geq 1}$. ∎
@@ 6.2. Double dual EPW surfaces as moduli spaces
Here we only elaborate on the case for a special Gushel–Mukai threefold $X$ over $S$ and the starting point is the Hilbert scheme $C(X)$ of conics on $X$. The surface $C(X)$ has two components $C_{\sigma}(X)$ and $C_{0}(X)$ such that $C_{\sigma}(X)\cong F(M_{S})\cong\mathbb{P}^{2}$ parameterizes the preimage of lines on $M_{S}$. The other component $C_{0}(X)$ is isomorphic to the blow up of the double dual EPW surface $\tilde{Y}^{\geq 2}_{A(X)^{\perp}}$ at the distinguished point $\tilde{\bm{p}}_{S}$ according to [18, Theorem 7.3]. These components intersect transversely along a smooth rational curve which is the exceptional curve on $C_{0}(X)$.
@@
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\rightarrow\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}_{2}}$
induced by the inflation functor is the natural double cover over $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}_{2}}=Y^{\geq 2}_{A(S)^{\perp}}$.
@@
The isomorphism $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ follows from [18, 28] as above. The identification of the fixed locus $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}_{2}}$ and the double cover is similar to [33, Section 6.5]. ∎
@@ 6.3. Double EPW surfaces as moduli spaces
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})\rightarrow\textsf{{M}}_{\sigma_{S}}(v_{2})_{\mathfrak{S}_{2}}$
induced by the inflation functor is the natural double cover over $\textsf{{M}}_{\sigma_{S}}(v_{2})_{\mathfrak{S}_{2}}=Y^{\geq 2}_{A(S)}$.
@@ 6.4. Description of the special period partners
$\mathscr{U}_{X}[2]\rightarrow G_{X}\rightarrow\mathscr{V}_{X}^{\vee}[1]\rightarrow$

### v2

@@ Abstract.
We show that the double dual EPW sextic and double EPW sextic associated with a strongly smooth Gushel–Mukai surface can be realized as moduli spaces of semistable objects with respect to a stability condition on the bounded derived category of it. Also, we observe that double dual EPW surface and double EPW surface associated with a special Gushel–Mukai threefold can be realized as moduli spaces of semistable objects on its Kuznetsov component. As an application, we refine a statement of Bayer and Perry about Gushel–Mukai threefolds with equivalent Kuznetsov components for the special ones.
@@ 1.3. Main results
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ of semistable objects with Mukai vector $v_{1}=(1,0,-1)$ is isomorphic to the double dual EPW sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ associated $S$.
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ of semistable objects with Mukai vector $v_{2}=(2,-H,2)$ is isomorphic to the double EPW sextics $\tilde{Y}^{\geq 1}_{A(S)}$ associated $S$.
@@ 5.4. The flopping walls and identification
It remains to show that $Y_{S}\cong X_{S}$ contracts to the moduli space $\textsf{{M}}_{\sigma_{S}}(1,0,-1)$ through the same symplectic resolution $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$ as in Theorem 5.1.
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ is isomorphic to $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$.
@@
where $Z\subset S$ is a length $2$ subscheme spanning a line not in $S$. Indeed, such a short exact sequence follows from the fact that the ideal sheaf $\mathscr{I}_{L/M_{S}}$ of every line $L\subset M_{S}$ can be written as the cokernel of a unique map $\mathscr{U}_{G}|_{M_{S}}\rightarrow\mathscr{V}_{G}|_{M_{S}}$, and each such map has the ideal sheaf of a line as the cokernel. One just takes the restriction onto $S\subset M_{S}$. The line spanned by $Z$ is not in $S$ because such a length two subscheme $Z$ has been already flopped in Proposition 5.13. So one contracts the Fano variety $F(M_{S})$ of lines, and therefore recovers $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$. ∎
Remark 5.15.
Moreover, the involution on $\textsf{{M}}_{\sigma_{S}}(v_{1})$ induced by the group action $\mathfrak{S}^{\vee}_{2}$ coincides with the covering involution on the double dual EPW sextics as the Hodge isometries of the corresponding birational involutions on $X_{S}$ are the same (see [33, Section 6.5] for the generic case). A point-wise description of the involution can be found in [18, Remark 7.8]. Also, the fixed locus $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}$ of the involution is exactly the ramification locus $Y^{\geq 2}_{A(S)^{\perp}}$.
@@ 6.2. Double dual EPW surfaces as moduli spaces
Here we only elaborate on the case for a special Gushel–Mukai threefold $X$ over $S$ and the starting point is the Hilbert scheme $C(X)$ of conics on $X$. The surface $C(X)$ has two components $C_{\sigma}(X)$ and $C_{0}(X)$ such that $C_{\sigma}(X)\cong F(M_{S})\cong\mathbb{P}^{2}$ parameterizes the preimage of lines on $M_{S}$. The other component $C_{0}(X)$ is isomorphic to the blow up of the double dual EPW surface associated with $X$ at the distinguished point $\tilde{\bm{p}}_{S}$ according to [18, Theorem 7.3]. These components intersect transversely along the exceptional curve on $C_{0}(X)$.
@@
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\rightarrow\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}$
induced by the inflation functor is the natural double cover over $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}=Y^{\geq 2}_{A(S)^{\perp}}$.
@@
The isomorphism $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ follows from [18, 28] as above. The identification of these branched double covers is similar to [33, Proposition 5.20]. One can do this because the involution on $C(X)$ induced by the branched cover $X\rightarrow M_{S}$ coincides with the one induced by the natural covering involution on $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ according to [18, Section 7.3]. In particular, the branched double cover for the double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ can be described in terms of conics on an open subset and therefore be identified. ∎
@@
However, in this case, one does not have enough information to give a double cover as the corresponding moduli space on the associated Gushel–Mukai fourfold is not clear.
@@ 6.3. Double EPW surfaces as moduli spaces
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})\rightarrow\textsf{{M}}_{\sigma_{S}}(v_{2})_{\mathfrak{S}^{\vee}_{2}}$
induced by the inflation functor is the natural double cover over $\textsf{{M}}_{\sigma_{S}}(v_{2})_{\mathfrak{S}^{\vee}_{2}}=Y^{\geq 2}_{A(S)}$.
@@ 6.4. Description of the special period partners
$\mathscr{U}_{X}[2]\rightarrow G_{X}\rightarrow\mathscr{V}_{X}^{\vee}[1]\rightarrow\mathscr{U}_{X}[3]$